OpenStudy (anonymous):

Rectangle A has a length of 2x + 6 and a width of 3x. Rectangle B has a length of x + 2 and an area of 12 square units greater than Rectangle A’s area. What is a simplified expression for the width of Rectangle B? x + 2 x + 1 6x + 6 6(x + 2)(x + 1)

OpenStudy (campbell_st):

area of rectangle A = (2x+6)(3x) = 6x^2 + 18x Reatangle B is 12 units larger so the area is 6x^2 + 18x + 12 = (x +2)(6x+6) reactangle B has an area of (x +2)(6x+6) with length (x+2) then the width is (6x + 6)

OpenStudy (experimentx):

(2x+6)3x = 6x^2 + 18x (x+2) (y) = 6x^2 + 18x + 12 (x+2) y = 6(x+1)(x+2) y = 6(x+1)

OpenStudy (anonymous):

@experimentX is the bottom one the answer?

OpenStudy (anonymous):

i'm sorry i mean the second one

OpenStudy (experimentx):

yes ... sorry I lost connection suddenly.

OpenStudy (anonymous):

ok